# Optimize Weight Rounding via Signed Gradient Descent for the Quantization of LLMs

### v4

@@ Appendix B View of distribution of tuned parameters
Appendix C Other results
We present the detailed accuracy results for 11 tasks using the LLaMA and Mistral models, ranging in size from 7B to 70B, at W2-W4 in Tables 10, 11, 12 and 13. The detailed perplexity (PPL) results are shown in Table 14. Overall, SignRound demonstrates a clear advantage in accuracy tasks, particularly in ultra-low bit quantization, achieving state-of-the-art performance compared to several popular weight quantization methods. In terms of perplexity (PPL), SignRound outperformed all other methods in 57 out of 112 scenarios, demonstrating its advantages. However, we observed that several quantization algorithms, including SignRound, exhibit sensitivity across different models and tasks. The reason for this sensitivity is detailed in Section 4.1.

### v5

@@ Appendix B View of distribution of tuned parameters
Table 14: Perplexity(PPL) ($\downarrow$) of Wikitext2, PTB, C4 and Wikitext tasks for LLaMA and Mistral models. we follow the source code of GPTQ for wikitext2, PTB and C4 PPL evaluation, while for wikitext, we adopt lm-eval-harness (Gao et al., 2023). NAN indicates not a number, while OOM denotes out of memory.
Appendix C More results
We present the detailed LLMs generalization results in Table 15, the accuracy is within 1% of the 16 bit benchmark after simple fine tuning on different types of models. The detailed accuracy results for 11 tasks using the LLaMA and Mistral models, ranging in size from 7B to 70B, at W2-W4 are provided in Tables 10, 11, 12 and 13. The detailed perplexity (PPL) results are shown in Table 14. Overall, SignRound demonstrates a clear advantage in accuracy tasks, particularly in ultra-low bit quantization, achieving state-of-the-art performance compared to several popular weight quantization methods. In terms of perplexity (PPL), SignRound outperformed all other methods in 83 out of 124 scenarios, demonstrating its advantages. However, we observed that several quantization algorithms, including SignRound, exhibit sensitivity across different models and tasks. The reason for this sensitivity is detailed in Section 4.1.
Gemma-2b BF16 32.87 63.44 52.73 65.04 76.71 22.03 29.80 69.27 64.26 74.20 40.19 53.69
Ours 32.97 63.07 51.59 65.43 76.12 22.03 30.00 69.39 63.90 73.53 39.33 53.40
Llama-2-7b-chat-hf FP16 46.40 71.05 57.80 66.38 76.39 30.23 33.40 79.76 69.68 73.82 44.20 59.01
Ours 45.45 70.37 57.06 66.14 76.33 30.35 32.60 80.64 72.92 73.36 43.52 58.97
Llama-3-8B-Instruct BF16 63.86 71.82 57.69 71.43 78.67 36.23 34.00 82.97 67.51 81.52 52.99 63.52
Ours 63.06 72.00 56.99 72.38 77.97 35.37 33.00 83.09 68.59 80.89 51.02 63.12
Mistral-7B-Instruct-v0.2 BF16 59.06 71.41 66.02 73.95 80.52 52.51 36.00 85.35 70.40 81.61 54.35 66.47
Ours 58.72 71.41 65.57 73.64 80.47 51.53 34.20 85.41 71.48 81.65 54.35 66.21
Mixtral-8x7B BF16 68.02 78.27 64.90 76.48 82.48 34.27 35.40 85.23 70.76 84.30 56.66 66.98
Ours 66.93 78.25 64.59 75.14 82.10 32.19 35.60 84.74 69.31 84.30 56.48 66.33
Mixtral-8x7B-Instruct BF16 68.85 77.18 67.67 76.87 83.51 49.69 36.80 88.50 71.84 86.99 62.20 70.00
Ours 68.24 77.90 67.45 77.19 83.35 48.84 37.20 87.83 70.04 87.12 62.29 69.77
Phi-3-mini-4k-instruct BF16 67.97 68.08 60.64 74.03 80.30 39.53 38.80 86.21 77.98 83.54 55.72 66.62
Ours 66.59 67.71 59.70 74.59 79.33 37.45 38.80 85.66 79.06 82.70 56.83 66.33
Table 15: The detail accuracies($\uparrow$) across 11 tasks(detailed in Section 4.1) with 1000 steps for LLMs at W4G128